# ECG-Chat: A Large ECG-Language Model for Cardiac Disease Diagnosis

### v1

@@ Abstract
The success of Multimodal Large Language Models (MLLMs) in the medical auxiliary field shows great potential, allowing patients to engage in conversations using physiological signal data. However, general MLLMs perform poorly in cardiac disease diagnosis, particularly in the integration of ECG data analysis and long-text medical report generation, mainly due to the complexity of ECG data analysis and the gap between text and ECG signal modalities. Additionally, models often exhibit severe stability deficiencies in long-text generation due to the lack of precise knowledge strongly related to user queries. To address these issues, we propose ECG-Chat, the first multitask MLLMs focused on ECG medical report generation, providing multimodal conversational capabilities based on cardiology knowledge. We propose a contrastive learning approach that integrates ECG waveform data with text reports, aligning ECG features with reports in a fine-grained manner. This method also results in an ECG encoder that excels in zero-shot report retrieval tasks. Additionally, expanding existing datasets, we constructed a 19k ECG diagnosis dataset and a 25k multi-turn dialogue dataset for training and fine-tuning ECG-Chat, which provides professional diagnostic and conversational capabilities. Furthermore, ECG-Chat can generate comprehensive ECG analysis reports through an automated LaTeX generation pipeline. We established a benchmark for the ECG report generation task and tested our model on multiple baselines. ECG-Chat achieved the best performance in classification, retrieval, multimodal dialogue, and medical report generation tasks. Our report template design has also been widely recognized by medical practitioners.
Figure 1: The ECG-Chat framework integrates several ECG-related medical text generation tasks within a single system. It includes: (1) medical question answering based on ECG and patient data (gray), (2) pathology classification (yellow), (3) diagnostic captioning (blue), (4) region-specific medical discussions (orange), and (5) comprehensive medical report generation (red). This makes ECG-Chat the first MLLM capable of generating diagnostic text by fully leveraging ECG signal data.
Introduction
In most regions of the world, doctors have absolute authority over the interpretation of pathological data for heart disease patients, creating an unequal dynamic that affects the trust between doctors and patients (Meyer and Ward 2008; Harnkham 2023). With the rapid development of MLLMs, intelligent and trustworthy interpretations based on medical report generation can effectively balance this inequality, enhancing patients’ willingness to seek medical attention and their confidence in doctors’ diagnoses. The pursuit of medical algorithm engineers is to leverage large models to generate accurate and rich personalized medical texts. This not only serves as an auxiliary tool to support doctors’ correct decision-making but also provides valuable medical advice to patient populations suffering from diseases, especially in developing countries where medical resources, particularly doctors, are scarce.
@@ Introduction
As a non-invasive physiological indicator detection method, electrocardiography (ECG) is a crucial tool for detecting early heart problems in patients. Some existing algorithms treat ECG data as a temporal physiological signal, focusing mainly on classification tasks (Chen et al. 2024c; Na et al. 2024; El-Ghaish and Eldele 2024). These models cannot help patients and can only serve as auxiliary tools for doctors. Also, Current ECG-language models (Liu et al. 2024b; Wan et al. 2024b) have not bridged the gap between ECG and text report modalities because most open ECG datasets lack comprehensive text descriptions. The structured, terminological, and highly repetitive phrase combinations poses challenge for the current Vision-Language models to migrate to the ECG field. Consequently, no model currently addresses both ECG report generation and question answering effectively. Additionally, large language models suffer from significant hallucinations in cardiology, making their accuracy unreliable (Zhou et al. 2024).
@@
In this paper, to address these challenges, we introduce ECG-Chat, a MLLM capable of generating long text reports for ECG. ECG-Chat uses the ECG encoder trained on the framework of contrastive learning. The original report is enhanced with waveform data, which shows excellent performance in retrieval tasks. After that, we used GPT-4o to build a dataset, ECG-Instruct, for instruction tuning ECG-LLM, which includes two forms: diagnosis and dialogue. With the ECG encoder, dataset and the LLM Vicuna-13B (Zheng et al. 2023), we constructed ECG-Chat that supports multiple functions such as report generation and ECG question and answer. At the same time, in order to solve the model’s hallucinations in report generation and medical knowledge, we built a prompt template for specific ECG diagnoses and a local knowledge base for retrieval enhancement generation (RAG). In addition, we built a pipeline that integrates patient information to generate detailed ECG reports. We tested the performance of the model on multiple tasks such as ECG-Report retrieval, ECG classification and ECG report generation. And established a benchmark for LLMs in ECG report generation.
In summary, the contributions of this paper are as follows:
@@
  • Waveform Data Enhancement: A contrastive learning method that combines ECG waveform data with text reports was proposed, aligning ECG features with report content at a fine-grained level. This resulted in the ECG encoder exhibiting the best performance in signal data representation.
@@
  • ECG-Instruct: We proposed a novel data generation pipeline, using existing datasets and GPT-4 to create an ECG instruction tuning dataset. The dataset contains 19K diagnosis and 25K dialogue datasets.
  • ECG-Chat: Utilizing our dataset, we fine-tuned Vicuna-13B to create an ECG-domain ECG-language model. Incremental learning based on LoRA (Hu et al. 2021) effectively prevented catastrophic forgetting of medical knowledge, and ECG signals were aligned with the LLM’s text embedding space. This enabled ECG-Chat to accept signal data, leverage Vicuna-13B’s conversational abilities, and extend them to medical dialogue and report generation tasks. We also proposed a Diagnosis-Driven Prompt (DDP) for ECG report generation, which effectively improved the accuracy of the model. Using an automated LaTeX pipeline, complex medical vocabulary was thoroughly explained and reported, presenting comprehensive patient medical history health advice in an easily understandable manner.
@@
Figure 2: The diagram depicts training process for ECG encoder (left part) and ECG-Chat (right part). The green font in the text report on the left is the waveform data enhancement.
Related Work
Cross-modal Medical Text Generation
@@
Cross-modal medical text generation refers to the process of leveraging multiple data modalities, such as images, text, and signals, to generate accurate and medically informed diagnostic or explanatory text, thereby supporting medical decision-making and treatment processes. Certain studies have successfully integrated multiple tasks into a single model, demonstrating remarkable functionality. (Zhang et al. 2023a; Xiong et al. 2023; Wu et al. 2023) As the field has progressed, unstructured data has been incorporated. (Wan et al. 2024a; Gupta et al. 2024; Yu et al. 2024; Chen et al. 2024a) Notably, image-text multimodal technology has shown exceptional performance in this domain, significantly enhancing diagnostic accuracy and medical efficiency. (Zhang et al. 2022; Huang et al. 2023; Lin et al. 2023; Zhang et al. 2022; Wu et al. 2023) However, research on signal-text multimodality, particularly in applications involving electrocardiograms, remains underexplored and holds substantial potential for future development.
Intelligent Interpretation of ECG
@@
The intelligent interpretation of ECG encompasses a range of tasks, including ECG retrieval, classification, and text generation, aimed at enhancing the accuracy and efficiency of cardiac diagnosis. In ECG retrieval, advanced algorithms are employed to search and identify relevant ECG patterns from large databases, enabling efficient comparisons and aiding in clinical decision-making. (Yu, Guo, and Sano 2024; Qiu et al. 2023b; Liu et al. 2024a) ECG classification has seen significant progress, with machine learning models being developed to automatically categorize ECG signals into various cardiac conditions, demonstrating high accuracy in identifying abnormalities. (Liu et al. 2024b; Markov et al. 2023; Qiu et al. 2023a; Xu et al. 2024; Davies, Monsen, and Mandic 2024; Hoang et al. 2024; Plagwitz et al. 2024; Huang et al. 2024) Despite these advancements, the generation of descriptive and diagnostic text from ECG data remains a relatively nascent area of research. (Li et al. 2024; Xu et al. 2024; Jiang et al. 2024; Steimetz et al. 2024; Yang et al. 2024; Wan et al. 2024b) Current studies focus on leveraging multimodal approaches to generate text that accurately reflects the underlying ECG signals, providing clinicians with clear and actionable insights. However, challenges such as ensuring the clinical relevance and interpretability of generated text, as well as addressing the nuances of complex cardiac conditions, remain areas requiring further exploration.
Method
Figure 3: Targeting the generation of long-form medical reports, ECG-Chat ensures the quality of medical text output through the integration of automated prompt tuning and the GraphRAG component. These enhanced methodologies are also applied across other chat functions within the system. As shown in (a), the Architecture of Diagnosis-Driven Prompts (indicated by the gray blocks) plays a critical role in report generation tasks. A classifier, implemented as a linear layer, is trained to estimate the probability of each potential diagnosis. The diagnosis with the highest probability is then selected as the final result, which is explicitly communicated to the LLMs using the keyword ”Present” in the prompt.
Architecture and General Pipeline
@@
ECG-Chat contains an end-to-end ECG diagnostic report generation pipeline. The ECG signals firstly pass through an ECG encoder, transforming the time series into feature representations. To enable the features fine-grained aligned with reports, we followed the contrastive learning architecture of the CoCa (Yu et al. 2022), and enhanced the text report with ECG waveform data. After that, similar to LLaVa-v1.5 (Liu et al. 2024c), we use a two-layer MLP adapter to align the feature space of the ECG encoder with the LLM through pretraining and fine-tuning. At the same time, we built a Dignosis-Driven Prompt based on linear classification for accurate ECG report generation, and built GraphRAG to address hallucinations in medical knowledge. Finally, a large language model of ECG diagnosis with the ability of report generation and multi-round conversations is constructed.
Aligning ECG Features with Text Reports
@@
In visual multimodal large language models, visual encoders are often obtained by training on large-scale image-text datasets by dual-encoder contrastiva learning(Radford et al. 2021; Li et al. 2023) or encoder-decoder captioning (Wang et al. 2022). CoCa (Yu et al. 2022) is a model that combines the two approches. It maps image to the same space as text representations by optimizing the contrastive loss through a dual-encoder. At the same time, the multimodal text decoder is used to optimize the captioning loss and improve the cross-modal generation ability of the model. Inspired by CoCa’s success in computer vision and ECG classification (Yu, Guo, and Sano 2024), we also use this model to extract features from ECG signals. Our architecture similarly uses contrastive loss and captioning loss, as shown in Fig. 2.
@@
$\mathcal{L}_{e2t}=\sum_{i}^{N}\log{\frac{\exp(x_{i}^{T}y_{i}/\sigma)}{{\textstyle\sum_{j=1}^{N}\exp(x_{i}^{T}y_{i}/\sigma)}}}$ (1)
$\mathcal{L}_{t2e}=\sum_{i}^{N}\log{\frac{\exp(y_{i}^{T}x_{i}/\sigma)}{{\textstyle\sum_{j=1}^{N}\exp(y_{i}^{T}x_{i}/\sigma)}}}$ (2)
@@
where $\mathcal{L}_{e2t}$ and $\mathcal{L}_{t2e}$ the contrastive loss of ecg-to-text and text-to-ecg, respectively. $x_{i}$ and $y_{i}$ are normalized embeddings of the ECG signal in the $i$-th pair and that of the text in $j$-th pair. $N$ is batch size, and $\sigma$ is the temperature to scale the logits. To address the insufficient amount of data, we use a pretrained frozen text encoder and fine-tune only its last few layers (Li et al. 2024).
@@
For the training data, we did not use diagnostic reports directly. Given the scarcity of datasets with free text reports, we appended corresponding ECG waveform data to the reports (Fig. 2). This approach artificially increases the distinction between samples, even when reports are identical, helping prevent contrastive loss from failing to converge during small-batch training. Additionally, incorporating waveform data ensures that the ECG encoding latent space captures more waveform information.
Multimodal Instruction Turning and Diagnosis-Driven Prompts for ECG
@@
The modality interface in ECG-Chat resembles LLaVA-v1.5 (Liu et al. 2024c). ECG encoding is embedded like text tokens and fed into large language models. Since ECG and text feature spaces differ, an adapter—a two-layer MLP—is used for conversion. Similar to LLaVA (Liu et al. 2023, 2024c), ECG-Chat undergoes two-stage pretraining: feature alignment and instruction fine-tuning, with the ECG encoder frozen in both stages. During feature alignment, only the linear projection layer is trained, while the LLM remains frozen. In the fine-tuning stage, the LLM is trained with LoRA (Hu et al. 2021). Due to the lack of existing datasets, we generated pretraining data and the ECG-Instruct dataset using GPT-4o, which will be detailed in the following chapter.
@@
Current large language models still face challenges in generating accurate medical reports, where precision is crucial. Inspired by the Diagnosis-Driven Prompt (DDP) approach used in radiology report generation (Jin et al. 2024; Chen et al. 2024d), we adapted DDP for ECG report generation, as illustrated in Fig. 3. Specifically, we use a linear layer to classify ECG feature vectors and obtain classification results. The label with the highest probability is embedded in the prompt as ”The label description is present.”. Given that ECG classification is a multi-label task, we further enhance accuracy by categorizing labels into disease, rhythm, and waveform. Labels with a predicted positive probability above a threshold are selected as classification results. Rhythm is treated as a single-label task, so the final output may include one or more labels. For instance, in Fig. 3, if the results are sinus rhythm (SR) and left bundle branch block (LBBB), the DDP ”Sinus rhythm is present; Left bundle branch block is present” is added to the model inputs.
Hallucination Elimination and Visualization Output
@@
As illustrated in Figure 3, our approach integrates three essential components to enhance the generation of ECG diagnostic reports. First, the GraphRAG component constructs a comprehensive knowledge graph from seven authoritative cardiology textbooks, enabling effective Retrieval-Augmented Generation (RAG) to mitigate hallucinations. This ensures that the generated text is grounded in established medical knowledge. Second, DSPy is employed for automated prompt tuning, which dynamically integrates patient data with retrieved knowledge, producing accurate and contextually relevant outputs. Finally, the LaTeX-based pipeline automates the creation of structured clinical reports, ensuring clarity and consistency in the final output. Detailed descriptions of these components are provided in Appendices A through C.
Dataset
Datasets for Contrastive Learning
@@
In the framework of ECG-Chat, the ECG encoder is trained using a large number of ECG-text pairs of data by contrastive learning. The existing open source datasets are relatively few, and the data size is also very small. Current multimodal models often integrate multiple datasets to improve generalization (Radford et al. 2021). Therefore, we also integrate three datasets for ECG encoder pre-training, namely MIMIC-IV-ECG (Gow et al. 2023), Champan-Shaoxing-Ningbo (CSN) (Zheng et al. 2020; Zheng, Guo, and Chu 2022) and Shandong Provincial Hospital (SPH) (Liu et al. 2022) datasets. The total size of the training dataset is 805K. For detailed data set description and preprocessing, please see the Appendix D.
@@
For the waveform data enhanced report, we used Python toolbox NeuroKit2 (Makowski et al. 2021) to extract the waveform information of each recording on lead II separately, including RR interval, PR interval, QRS complex duration, QT/QTc interval, and the peaks of P, R and T waves. The waveform data is added directly after the text report, as shown in the Fig. 2.
@@
Figure 4: An example of the ECG-Instruct dataset. The context is used as the prompt of GPT-4 to describe an ECG. The bottom half is the responses generated by GPT-4 for building our instruction-following dataset.
Datasets for Instrution Tuning

### v2

@@ Abstract
The success of Multimodal Large Language Models (MLLMs) in the medical auxiliary field shows great potential, allowing patients to engage in conversations using physiological signal data. However, general MLLMs perform poorly in cardiac disease diagnosis, particularly in the integration of ECG data analysis and medical report generation, mainly due to the complexity of ECG data analysis and the gap between text and ECG signal modalities. To address these issues, we propose ECG-Chat, a multitask MLLMs focused on ECG medical report generation, providing multimodal conversational capabilities based on cardiology knowledge. We propose a contrastive learning approach that integrates ECG waveform data with text reports, aligning ECG features with reports in a fine-grained manner. This method also results in an ECG encoder that excels in zero-shot report retrieval tasks. Additionally, expanding existing datasets, we constructed a 19k ECG diagnosis dataset and a 25k multi-turn dialogue dataset for training and fine-tuning ECG-Chat, which provides professional diagnostic and conversational capabilities. Furthermore, ECG-Chat can generate comprehensive ECG analysis reports through an automated LaTeX generation pipeline. We established a benchmark for the ECG report generation task and tested our model on multiple baselines. ECG-Chat achieved the best performance in classification, retrieval, and medical report generation tasks. Our code is available at https://github.com/YubaoZhao/ECG-Chat.
Index Terms:
ECG, Large Language Model, Multimodal
I Introduction
@@
With the rapid development of multimodal large language models (MLLMs), these advanced systems show great promise in assisting patients by integrating diverse modalities such as text, images, and medical data to enhance healthcare delivery and decision-making. As a non-invasive physiological indicator detection method, electrocardiography (ECG) is a crucial tool for detecting early heart problems in patients. Some existing algorithms treat ECG data as a temporal physiological signal, focusing mainly on classification tasks [1, 2, 3, 4]. These models cannot help patients and can only serve as auxiliary tools for doctors. Also, Current ECG-language models [5, 6] have not bridged the gap between ECG and text report modalities because most open ECG datasets lack comprehensive text descriptions. The structured, terminological, and highly repetitive phrase combinations poses challenge for the current Vision-Language models to migrate to the ECG field. Consequently, no model currently addresses both ECG report generation and question answering effectively. Additionally, large language models suffer from significant hallucinations in cardiology, making their accuracy unreliable.
@@
In this paper, to address these challenges, we introduce ECG-Chat, a MLLM capable of generating long text reports for ECG. ECG-Chat uses the ECG encoder trained on the framework of contrastive learning. The original report is enhanced with waveform data, which shows excellent performance in retrieval tasks. After that, we used GPT-4o to build a dataset, ECG-Instruct, for instruction tuning ECG-LLM, which includes two forms: diagnosis and dialogue. With the ECG encoder, dataset and the LLM Vicuna-13B [7], we constructed ECG-Chat that supports multiple functions such as report generation and ECG question and answer. At the same time, in order to solve the model’s hallucinations in report generation and medical knowledge, we built a prompt template for specific ECG diagnoses and a local knowledge base for retrieval enhancement generation (RAG). In addition, we built a pipeline that integrates patient information to generate detailed ECG reports. We tested the performance of the model on multiple tasks such as ECG-Report retrieval, ECG classification and ECG report generation. And established a benchmark for LLMs in ECG report generation.
In summary, this paper makes the following contributions:
@@
  • Waveform Data Enhancement: Introduced a contrastive learning method aligning ECG waveform data with text reports, achieving state-of-the-art ECG encoder performance.
@@
  • ECG-Instruct: Developed a novel pipeline using GPT-4 to generate a 19K-diagnosis and 25K-dialogue ECG instruction tuning dataset.
  • ECG-Chat: Fine-tuned Vicuna-13B with LoRA to prevent catastrophic forgetting, aligning ECG signals with LLM text embeddings. Proposed a Diagnosis-Driven Prompt (DDP) for accurate ECG report generation, incorporating automated LaTeX-based medical vocabulary explanations for comprehensive and accessible reports.
@@
Figure 1: The diagram depicts training process for ECG encoder (left part) and ECG-Chat (right part).
II Related Work
II-A Cross-modal Medical Text Generation
@@
Cross-modal medical text generation refers to the process of leveraging multiple data modalities, such as images, text, and signals, to generate accurate and medically informed diagnostic or explanatory text, thereby supporting medical decision-making and treatment processes. Certain studies have successfully integrated multiple tasks into a single model, demonstrating remarkable functionality [8]. As the field has progressed, unstructured data has been incorporated. [9, 10, 11, 12] Notably, image-text multimodal technology has shown exceptional performance in this domain, significantly enhancing diagnostic accuracy and medical efficiency. [13] However, research on signal-text multimodality, particularly in applications involving electrocardiograms, remains underexplored and holds substantial potential for future development.
II-B Intelligent Interpretation of ECG
@@
The intelligent interpretation of ECG encompasses a range of tasks, including ECG retrieval, classification, and text generation, aimed at enhancing the accuracy and efficiency of cardiac diagnosis. In ECG retrieval, advanced algorithms are employed to search and identify relevant ECG patterns from large databases, enabling efficient comparisons and aiding in clinical decision-making[14]. ECG classification has seen significant progress, with machine learning models being developed to automatically categorize ECG signals into various cardiac conditions, demonstrating high accuracy in identifying abnormalities[15, 5, 1, 2, 3, 4, 16, 17]. Despite these advancements, the generation of descriptive and diagnostic text from ECG data remains a relatively nascent area of research [2, 18, 6]. Current studies focus on leveraging multimodal approaches to generate text that accurately reflects the underlying ECG signals, providing clinicians with clear and actionable insights. However, challenges such as ensuring the clinical relevance and interpretability of generated text, as well as addressing the nuances of complex cardiac conditions, remain areas requiring further exploration.
III Method
Figure 2: ECG-Chat improves long-form medical report generation by integrating automated prompt tuning and GraphRAG into its system. This approach extends to other chat functions as well. As illustrated in (a), a key component is a classifier (shown as gray blocks), built as a linear layer, trained to predict the likelihood of various diagnoses.
III-A Architecture and General Pipeline
@@
ECG-Chat contains an end-to-end ECG diagnostic report generation pipeline. The ECG signals firstly pass through an ECG encoder, transforming the time series into feature representations. To enable the features fine-grained aligned with reports, we followed the contrastive learning architecture of the CoCa [19], and enhanced the text report with ECG waveform data. After that, similar to LLaVa-v1.5 [20], we use a two-layer MLP adapter to align the feature space of the ECG encoder with the LLM through pretraining and fine-tuning. At the same time, we built a Dignosis-Driven Prompt based on linear classification for accurate ECG report generation, and built GraphRAG to address hallucinations in medical knowledge. Finally, a large language model of ECG diagnosis with the ability of report generation and multi-round conversations is constructed.
III-B Aligning ECG Features with Text Reports
@@
In visual multimodal large language models, visual encoders are often obtained by training on large-scale image-text datasets by dual-encoder contrastiva learning[21] or encoder-decoder captioning [22]. CoCa [19] is a model that combines the two approches. It maps image to the same space as text representations by optimizing the contrastive loss through a dual-encoder. At the same time, the multimodal text decoder is used to optimize the captioning loss and improve the cross-modal generation ability of the model. Inspired by CoCa’s success in computer vision and ECG classification [15], we also use this model to extract features from ECG signals. Our architecture similarly uses contrastive loss and captioning loss, as shown in Fig. 1.
@@
$\mathcal{L}_{e2t}$ $=\sum_{i}^{N}\log{\frac{\exp(x_{i}^{T}y_{i}/\sigma)}{{\textstyle\sum_{j=1}^{N}\exp(x_{i}^{T}y_{j}/\sigma)}}}$ (1)
$\mathcal{L}_{t2e}$ $=\sum_{i}^{N}\log{\frac{\exp(y_{i}^{T}x_{i}/\sigma)}{{\textstyle\sum_{j=1}^{N}\exp(y_{i}^{T}x_{j}/\sigma)}}}$ (2)
@@
where $\mathcal{L}_{e2t}$ and $\mathcal{L}_{t2e}$ the contrastive loss of ecg-to-text and text-to-ecg, respectively. $x_{i}$ and $y_{i}$ are normalized embeddings of the ECG signal in the $i$-th pair and that of the text in $j$-th pair. $N$ is batch size, and $\sigma$ is the temperature to scale the logits. To speed up the training, we use a pretrained text encoder and fine-tune only its last few layers [23].
@@
For the training data, we did not use diagnostic reports directly. Given the scarcity of datasets with text reports, we appended corresponding waveform data to the reports (Fig. 1). This approach artificially increases the distinction between samples, even when reports are identical, helping prevent contrastive loss from failing to converge during small-batch training. Additionally, incorporating waveform data ensures that the ECG encoding latent space captures more waveform information.
III-C Multimodal Instruction Turning and Inference
@@
The modality interface in ECG-Chat resembles LLaVA-v1.5 [20]. ECG encoding is embedded like text tokens and fed into LLMs. Since ECG and text feature spaces differ, an adapter—a two-layer MLP—is used for conversion. ECG-Chat undergoes two-stage pretraining: feature alignment and instruction fine-tuning, with the ECG encoder frozen in both stages. During feature alignment, only the linear projection layer is trained, while the LLM remains frozen. In the fine-tuning stage, the LLM is trained with LoRA [24]. Due to the lack of existing datasets, we generated pretraining data and the ECG-Instruct dataset using GPT-4o, which will be detailed in the following chapter.
@@
Large language models struggle with generating precise medical reports. Inspired by the Diagnosis-Driven Prompt (DDP) method in radiology [25, 26], we adapted DDP for ECG report generation (Fig. 2). A linear layer classifies ECG feature vectors, embedding the most probable label in the prompt (e.g., ”The label description is present.”). For multi-label ECG classification, we enhance accuracy by grouping labels into disease, rhythm, and waveform categories. Labels with probabilities above a threshold are included, treating rhythm as single-label. For example, in Fig. 2, outputs like sinus rhythm (SR) and left bundle branch block (LBBB) are embedded as ”Sinus rhythm is present; Left bundle branch block is present” in the model input.
III-D Hallucination Elimination and Visualization Output
@@
Our approach integrates three essential components to enhance the generation of ECG diagnostic reports. First, the GraphRAG component constructs a comprehensive knowledge graph from seven authoritative cardiology textbooks, enabling effective Retrieval-Augmented Generation (RAG) to mitigate hallucinations. This ensures that the generated text is grounded in established medical knowledge. Second, DSPy is employed for automated prompt tuning, which dynamically integrates patient data with retrieved knowledge, producing accurate and contextually relevant outputs. Finally, the LaTeX-based pipeline automates the creation of structured clinical reports, ensuring clarity and consistency in the final output. Detailed descriptions of these components are provided in Appendices A through C.
IV Dataset
IV-A Datasets for Contrastive Learning
@@
In the framework of ECG-Chat, the ECG encoder is trained using a large number of ECG-text pairs of data by contrastive learning. The existing open source datasets are relatively few, and the data size is also very small. Current multimodal models often integrate multiple datasets to improve generalization [21]. Therefore, we also integrate three datasets for ECG encoder pre-training, namely MIMIC-IV-ECG [27], Champan-Shaoxing-Ningbo (CSN) [28, 29] and Shandong Provincial Hospital (SPH) [30] datasets. The total size of the training dataset is 805K. For detailed data set description and preprocessing, please see the Appendix D.
@@
For the waveform data enhanced report, we used Python toolbox NeuroKit2 [31] to extract the waveform information of each recording on lead II separately, including RR interval, PR interval, QRS complex duration, QT/QTc interval, and the peaks of P, R and T waves. The waveform data is added directly after the text report, as shown in the Fig. 1.
@@
Figure 3: An example of the ECG-Instruct dataset. The context is used as the prompt of GPT-4 to describe an ECG.
IV-B Datasets for Instrution Tuning